an air conditioner removes heat steadily form a house at a rate of 850 kj/min while drawing electric power at a rate of 6 kW. determine the COP of this air conditioner


1
Expert's answer
2021-01-20T10:24:59-0500

The coefficient of performance of the air-conditioner can be calculated as follows:


"COP_R=\dfrac{Q_L}{W_{net, in}},"

here, "Q_L"is the total heat removed by the air conditioner, "W_{net, in}"is the power input into the air conditioner.

Then, we get:


"COP_R=\dfrac{850\ \dfrac{kJ}{min}\cdot\dfrac{1000\ J}{1\ kJ}\cdot\dfrac{1\ min}{60\ s}}{6\ kW\cdot\dfrac{1000\ W}{1\ kW}}=2.36"

Answer:

"COP_R=2.36"